OpenStudy (anonymous):

Solve the problems! f+16=31 Solve for letter! You answer and we see if we get the same answer!

OpenStudy (ash2326):

f+16=31 f=31-16 f=15

OpenStudy (hoblos):

f=15

OpenStudy (anonymous):

That's what I got.

OpenStudy (turingtest):

if you already have think you have the answer post what you think it is please!

OpenStudy (anonymous):

f=15

OpenStudy (turingtest):

that saves us some explaining!

OpenStudy (anonymous):

k